# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ II.3 SymTFT Description of Gauging
The preceding discussion focused on how the SymTFT construction can be used to identify distinct gapped phases of a $1+1$d $G$-symmetric system. In fact, these methods have further utility in that the SymTFT framework can be employed to relate the universal aspects of $1+1$d theories which may be invariant under different symmetry groups but are nonetheless related via gauging. Consequently, dualities which arise from gauging–such as the Kramers-Wannier duality [4]–can also be incorporated within this perspective.
Stated more explicitly, dualities of a $1+1$d gapped $G$-symmetric phase are realized by changing the gapped boundary condition at the reference boundary [32, 37]. Consider a $G$-symmetric $1+1$d gapped phase and let $H$ be a subgroup of $G$ that we wish to gauge, with $\mathcal{L}_{H}$ the Lagrangian subgroup that is condensed on the dynamical boundary in the SymTFT to realize SSB to $H$. In the slab picture, gauging $H$ is realized by changing the reference boundary condition to the $\mathcal{L}_{H}$ condensed boundary. If $\mathcal{L}^{\prime}_{H}$ is a different Lagrangian subgroup that realizes SSB to $H$ (i.e., these differ by an $H$-SPT), then changing the reference boundary to $\mathcal{L}^{\prime}_{H}$ in the SymTFT implements twisted gauging in the $1+1$d theory.
This prescription allows us to straightforwardly identify the kinematic data of the new symmetry. As an example, let us take $G=\mathbb{Z}_{4}$ and suppose that $H=\mathbb{Z}_{2}$ is the subgroup we wish to gauge. By Eq. (8), the reference boundary is specified by the choice $\mathcal{L}=\langle e^{2},m^{2}\rangle$. The symmetry operators are the confined line operators at the reference boundary, generated by $e$ and $m$. Since $e^{2}$ and $m^{2}$ are condensed at the reference boundary, $e$ and $m$ generate a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Moreover, the nontrivial braiding between $e$ and $m$ corresponds to a mixed ’t Hooft anomaly between the two $\mathbb{Z}_{2}$ symmetries at the boundary. Note that this braiding cannot be removed by simply attaching line operators corresponding to condensed anyons at the reference boundary to the line operators for $e$ and $m$ [91, 33]. This mixed ’t Hooft anomaly enforces the constraint that there does not exist a short-range entangled, gapped ground state that is invariant under the global $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Indeed, any pattern of anyon condensation on the dynamical boundary must share at least one nontrivial anyon with the reference boundary, i.e., there does not exist any magnetic Lagrangian subgroup on the dynamical boundary [92]. This implies that any gapped dynamical boundary must partially break the symmetry; this is the SymTFT restatement of the familiar fact that there is no trivial symmetry-preserving gapped state for the anomalous $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry in $1+1$d.
@@ III Review: Partial Anyon Condensation and Gapless SPTs
While one naturally expects the SymTFT to accurately encapsulate all $G$-symmetric gapped phases (given the bulk-boundary correspondence), it has recently become clear that the utility even extends to certain families of gapless phases of matter [36, 37, 39, 38, 42, 43, 44, 45]. Indeed, when anyon condensation at the dynamical boundary is not fully confining (equivalently, the condensable algebra is not maximal), the dynamical boundary describes the universal kinematic properties of a gapless $1+1$ $G$-symmetric phase. In this Section, we review this connection between partial anyon condensation and gapless $1+1$d phases, focusing in particular on the description of intrinsically gapless SPTs (igSPTs) proposed in Ref. [36].
In order to set the context for igSPTs and partial anyon condensation, let us consider a $1+1$d lattice spin system, which has a tensor product structure generated by finite-dimensional local Hilbert spaces at each site, and an on-site symmetry $G$. As shown recently, this is equivalent to the statement that $G$ is non-anomalous (so that $H^{3}(G,U(1))=\mathbb{Z}_{1}$[93]. At low energies, it may so happen that the full symmetry group does not act faithfully; specifically, there may be a normal subgroup $A\lhd G$ which only acts nontrivially on some gapped degrees of freedom, such that the effective symmetry group of the low-energy theory is given by the quotient group $G_{low}=G/A$. This structure is described by the short exact sequence,
@@
The key idea is that while $G$ is on-site and therefore non-anomalous, the action of $G_{low}$ on the low-energy degrees of freedom can be anomalous, with an anomaly captured by a cohomology class $[\omega]\in H^{3}(G_{low},U(1))$.
Since $G$ is an on-site symmetry, this emergent anomaly of $G_{low}$ is trivialized by the symmetry extension to $G$ [54]. In this way, the low energy physics of the $1+1$d $G$-symmetric lattice spin system can mimic the behavior of the boundary of a $2+1$d $G_{low}$-SPT, the anomaly of which enforces either an SSB or a gapless, symmetric phase. By extending $G_{low}$ to $G$ by including gapped degrees of freedom transforming nontrivially under $A$, it is hence possible to realize a protected gapless phase whose robustness is guaranteed by the emergent $G_{low}$ anomaly. Such anomalous gapless phases are referred to as intrinsically gapless SPTs (igSPTs). Despite being gapless, igSPTs are characterized by nontrivial string order parameters and support topologically protected edge modes [54], justifying the appellation “topological”. The “intrinsically gapless” nature of igSPTs refers to the fact that there do not exist purely gapped SPT analogues of these states in $1+1$d. For instance, as we review below, there exists a $1+1$d igSPT protected by $\mathbb{Z}_{4}$ symmetry which, on its own, is not sufficient to protect any gapped SPTs in $1+1$d (since $H^{2}(\mathbb{Z}_{4},U(1))=\mathbb{Z}_{1}$).
Although a priori one may not expect the topological data defining a quantum double $\mathcal{D}(G)$ to be sufficient to fully constrain the dynamics of a $G$-symmetric gapless theory, it turns out that certain boundaries of the former do encode the universal, kinematic data of the emergent anomaly discussed above. However, this requires relaxing the gapped boundary condition enforced on the dynamical boundary in Section II and considering gapless boundaries of $\mathcal{D}(G)$, which permit (partial) characterization in terms of the algebraic data governing anyon condensation. Lifting the gapped boundary condition and allowing gapless edges of the bulk TO is achieved by relaxing the constraint that the corresponding subalgebras of condensed bulk anyons $\mathcal{A}$ are Lagrangian. As illustrated in Fig. 1(a), condensing a condensable but non-Lagrangian subalgebra of anyons $\mathcal{A}$ within a subregion of the bulk TO leads to a new TO $\mathcal{D}(G)/\mathcal{A}$ within the subregion, with gapped domain walls separating the two distinct anyon theories. On the other hand, condensing a non-Lagrangian $\mathcal{A}$ at a boundary leads to a gapless boundary theory (see Fig. 1(c)). As we will review below, the reduced $\mathcal{D}(G)/\mathcal{A}$ TO encodes the effective low-energy symmetry group $G_{low}$ of the corresponding $1+1$d gapless theory.
@@ III.1 Classification of Condensable Subalgebras and Partial Anyon Condensation
In Sec. II, we reviewed the correspondence between gapped boundaries of $\mathcal{D}(G)$ topological orders, Lagrangian subalgebras, and $1+1$d $G$-symmetric gapped phases. Here, we review the classification of partial anyon condensation in $2+1$d anyon theories described by the quantum double $\mathcal{D}(G)$; such partial condensations mathematically correspond to condensable (but not Lagrangian) subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$, i.e., sets of bosonic anyons which all braid trivially with each other. Relaxing the Lagrangian condition amounts to lifting the condition that every anyon $b\notin\mathcal{A}$ braid nontrivially with at least one anyon $a\in\mathcal{A}$. In this scenario, anyons which lie outside $\mathcal{A}$ but braid trivially with anyons in $\mathcal{A}$ will remain deconfined anyons after $\mathcal{A}$ is condensed.
@@
The holographic dictionary described in Ref. [36] showed how this data can be understood as describing the symmetry fractionalization of $A$ and $H$ on the boundary. In the case when $H=G$, the condensable algebra $\mathcal{A}$ contains no pure charges, and the resulting boundary does not break any symmetry; since we are interested in $G$-symmetric $1+1$d phases, we restrict to this case for the remainder of this Section. Note that the condensable algebra is Lagrangian if and only if $A=H$. In this case, $\epsilon$ is fully determined by $\eta$, and the classification reduces to the case discussed in Sec. II.
@@
where $a$ labels the condensed anyons and the integer $n_{a}\in\mathbb{Z}_{>0}$ labels the corresponding fusion multiplicity. However, this decomposition does not fully characterize the condensable algebra [83, 84]. Nevertheless, this decomposition into condensed anyons is useful in developing an intuitive picture for the condensation. Since the precise relation between the data $A\subset H$, $\epsilon$, and $\eta$ with the set of condensed anyons is somewhat involved for general $G$, we defer a detailed discussion to Appendix A and discuss the simpler case when $G$ is Abelian here.
@@
In this case, the data encoded in $\epsilon$ and $\eta$ can be repackaged into a (not necessarily Lagrangian) subgroup of condensed anyons. First, note that the condition Eq. (21b) shows that $\epsilon(a,\cdot):H\to U(1)$ is a $1$d irrep of $H$. This gives the charge of the condensed dyon with flux $a$: for every flux $a\in A$, the dyon labeled by $(a,\epsilon(a,\cdot))$ is condensed. The condition Eq. (21a) ensures that the condensed anyons form a group: the charge bound to the flux $ab$ is equal to the product of the charges associated to $a$ and $b$. To understand the final condition Eq. (21c), note that since $\eta$ is a 2-cocycle on $A$, the ratio $\Omega(a,b)=\eta(b,a)/\eta(a,b)$ satisfies the same properties in Eq. (4) as the function $\Omega$ defined in Eq. (3). As before, $\Omega(a,\cdot)$ can be interpreted as the $A$-charge associated to the flux $a$. Thus, the final condition is a consistency condition that states that the restriction of the $H$-charge $\epsilon(a,\cdot)$ agrees with the $A$-charge $\Omega(a,\cdot)$ when $H$ is restricted to $A$. Finally, we note that the third condition Eq. (21c) implies that
@@ III.2 SymTFT for igSPTs
The key insight of Ref. [36] was that the data classifying condensable subalgebras of a quantum double $\mathcal{D}(G)$ is in one-to-one correspondence with the data characterizing the symmetry fractionalization patterns of $G$-symmetric igSPTs in $1+1$d. Here, we briefly sketch out this correspondence and refer the reader to Ref. [36] for a more detailed exposition.
Consider a $G$-symmetric $1+1$d spin system. Let us assume that, at low energies, a normal subgroup $A\lhd G$ acts trivially, such that the effective low-energy symmetry group is given by the quotient $G_{low}=G/A$. Now, on a system with periodic boundary conditions, we expect that the symmetry operator $U_{a}$, with $a\in A$, acts trivially. However, since $A$ only acts nontrivially on the gapped degrees of freedom, the symmetry may fractionalize on a system with open boundary conditions, such that $U_{a}=U_{a}^{L}\otimes U_{a}^{R}$, where $U_{a}^{L/R}$ has spatial support only on the left/right edge of the system. Since the symmetry operators $U_{a}$ form a linear representation of $A$, the restricted symmetry operators $U_{a}^{\alpha=L/R}$ need only form a projective representation:
@@
where $\eta\in\mathcal{Z}(A,U(1))$ is a 2-cocycle. Physically, the presence of a nontrivial $[\eta]$ implies that the gapped degrees of freedom form an $A$-SPT.
@@
where $\epsilon(a,g)\in U(1)$ is a phase that may be interpreted as measuring the fractionalized charge under $G$ that $U^{\alpha}_{a}$ carries. As shown in Ref. [36], the 3-cocycle $[\alpha]\in H^{3}(A,U(1))$ characterizing the emergent anomaly of the igSPT can be extracted from the pair of functions $(\eta,\epsilon)$.
@@ III.3 Example: $\mathbb{Z}_{4}$ igSPT
To illustrate the main idea behind the SymTFT perspective on igSPTs, let us consider in detail the example of the $1+1$d $\mathbb{Z}_{4}$ gapless SPT first introduced in Ref. [54]. The $\mathbb{Z}_{4}$ gapless SPT is characterized by the short exact sequence
@@
where the low-energy symmetry $G_{low}=\mathbb{Z}_{2}$ exhibits an emergent anomaly characterized by the nontrivial 3-cocycle class $\alpha\in H^{3}(\mathbb{Z}_{2},U(1))\cong\mathbb{Z}_{2}$, which is trivialized by the extension to the full symmetry group $G=\mathbb{Z}_{4}$ [54]. The low energy behavior of the $\mathbb{Z}_{4}$ gapless SPT is identical to that of the symmetry-preserving boundary of the $2+1$d $\mathbb{Z}_{2}$ Levin-Gu SPT [94], which carries the same anomaly $\alpha$. The physical consequence of this anomaly is that it prevents a $1+1$d theory from realizing a short-range entangled state that is symmetry preserving, thereby enforcing either gaplessness or spontaneous symmetry breaking.
@@
That the data encapsulating the universal, low-energy behavior of the $\mathbb{Z}_{4}$ gapless SPT and the boundary of the Levin-Gu SPT are identical suggests that the holographic dual for the former should be the gauged version of the Levin-Gu SPT, which is given by the doubled semion model TO $\mathcal{D}(\mathbb{Z}^{\alpha}_{2})$. On the other hand, since the global symmetry for the gapless SPT is $\mathbb{Z}_{4}$, one might expect that the corresponding bulk anyon theory is given by the $\mathcal{D}(\mathbb{Z}_{4})$ TO. As it turns out, the resolution to this ostensible puzzle is that both are true since the doubled semion TO can be obtained via the $\mathcal{D}(\mathbb{Z}_{4})$ anyon theory by condensing the $e^{2}m^{2}$ anyon in the latter. The physical picture is provided by Fig. 3: the blue region denotes the $\mathcal{D}(\mathbb{Z}_{4})$ TO while the green region denotes the doubled semion TO, obtained via condensation of $e^{2}m^{2}$ in the former. Thus, the condensable algebra generated by $\langle e^{2}m^{2}\rangle$ in $\mathcal{D}(\mathbb{Z}_{4})$ precisely encodes the local low-energy properties of a $1+1$d $\mathbb{Z}_{4}$ symmetric gapless theory (red dotted line in Fig. 3), which correspond to a $\mathbb{Z}_{2}$ symmetry along with the Levin-Gu anomaly $\alpha$. As discussed above, this is precisely the data that characterizes the $\mathbb{Z}_{4}$ igSPT.

### v2

@@ II.3 SymTFT Description of Gauging
The preceding discussion focused on how the SymTFT construction can be used to identify distinct gapped phases of a 1+11+1d $G$-symmetric system. In fact, these methods have further utility in that the SymTFT framework can be employed to relate the universal aspects of 1+11+1d theories which may be invariant under different symmetry groups but are nonetheless related via gauging. Consequently, dualities which arise from gauging–such as the Kramers-Wannier duality [4]–can also be incorporated within this perspective.
Stated more explicitly, dualities of a 1+11+1d gapped $G$-symmetric phase are realized by changing the gapped boundary condition at the reference boundary [33, 38]. Consider a $G$-symmetric 1+11+1d gapped phase and let $H$ be a subgroup of $G$ that we wish to gauge, with $\mathcal{L}_{H}$ the Lagrangian subgroup that is condensed on the dynamical boundary in the SymTFT to realize SSB to $H$. In the slab picture, gauging $H$ is realized by changing the reference boundary condition to the $\mathcal{L}_{H}$ condensed boundary. If $\mathcal{L}^{\prime}_{H}$ is a different Lagrangian subgroup that realizes SSB to $H$ (i.e., these differ by an $H$-SPT), then changing the reference boundary to $\mathcal{L}^{\prime}_{H}$ in the SymTFT implements twisted gauging in the 1+11+1d theory.
This prescription allows us to straightforwardly identify the kinematic data of the new symmetry. As an example, let us take $G=\mathbb{Z}_{4}$ and suppose that $H=\mathbb{Z}_{2}$ is the subgroup we wish to gauge. By Eq. (8), the reference boundary is specified by the choice $\mathcal{L}=\langle e^{2},m^{2}\rangle$. The symmetry operators are the confined line operators at the reference boundary, generated by $e$ and $m$. Since $e^{2}$ and $m^{2}$ are condensed at the reference boundary, $e$ and $m$ generate a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Moreover, the nontrivial braiding between $e$ and $m$ corresponds to a mixed ’t Hooft anomaly between the two $\mathbb{Z}_{2}$ symmetries at the boundary. Note that this braiding cannot be removed by simply attaching line operators corresponding to condensed anyons at the reference boundary to the line operators for $e$ and $m$ [96, 34]. This mixed ’t Hooft anomaly enforces the constraint that there does not exist a short-range entangled, gapped ground state that is invariant under the global $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Indeed, any pattern of anyon condensation on the dynamical boundary must share at least one nontrivial anyon with the reference boundary, i.e., there does not exist any magnetic Lagrangian subgroup on the dynamical boundary [97]. This implies that any gapped dynamical boundary must partially break the symmetry; this is the SymTFT restatement of the familiar fact that there is no trivial symmetry-preserving gapped state for the anomalous $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry in 1+11+1d.
@@ III Review: Partial Anyon Condensation and Gapless SPTs
While one naturally expects the SymTFT to accurately encapsulate all $G$-symmetric gapped phases (given the bulk-boundary correspondence), it has recently become clear that the utility even extends to certain families of gapless phases of matter [37, 38, 40, 39, 43, 44, 45, 46]. Indeed, when anyon condensation at the dynamical boundary is not fully confining (equivalently, the condensable algebra is not maximal), the dynamical boundary describes the universal kinematic properties of a gapless $1+1$ $G$-symmetric phase. In this Section, we review this connection between partial anyon condensation and gapless 1+11+1d phases, focusing in particular on the description of intrinsically gapless SPTs (igSPTs) proposed in Ref. [37].
In order to set the context for igSPTs and partial anyon condensation, let us consider a 1+11+1d lattice spin system, which has a tensor product structure generated by finite-dimensional local Hilbert spaces at each site, and an on-site symmetry $G$. As shown recently, this is equivalent to the statement that $G$ is non-anomalous (so that $H^{3}(G,U(1))=\mathbb{Z}_{1}$) [98]. At low energies, it may so happen that the full symmetry group does not act faithfully; specifically, there may be a normal subgroup $A\lhd G$ which only acts nontrivially on some gapped degrees of freedom, such that the effective symmetry group of the low-energy theory is given by the quotient group $G_{low}=G/A$. This structure is described by the short exact sequence,
@@
The key idea is that while $G$ is on-site and therefore non-anomalous, the action of $G_{low}$ on the low-energy degrees of freedom can be anomalous, with an anomaly captured by a cohomology class $\omega]\in H^{3}(G_{low},U(1))$.
Since $G$ is an on-site symmetry, this emergent anomaly of $G_{low}$ is trivialized by the symmetry extension to $G$ [57]. In this way, the low energy physics of the 1+11+1d $G$-symmetric lattice spin system can mimic the behavior of the boundary of a 2+12+1d $G_{low}$-SPT, the anomaly of which enforces either an SSB or a gapless, symmetric phase. By extending $G_{low}$ to $G$ by including gapped degrees of freedom transforming nontrivially under $A$, it is hence possible to realize a protected gapless phase whose robustness is guaranteed by the emergent $G_{low}$ anomaly. Such anomalous gapless phases are referred to as intrinsically gapless SPTs (igSPTs). Despite being gapless, igSPTs are characterized by nontrivial string order parameters and support topologically protected edge modes [57], justifying the appellation “topological”. The “intrinsically gapless” nature of igSPTs refers to the fact that there do not exist purely gapped SPT analogues of these states in 1+11+1d. For instance, as we review below, there exists a 1+11+1d igSPT protected by $\mathbb{Z}_{4}$ symmetry which, on its own, is not sufficient to protect any gapped SPTs in 1+11+1d (since $H^{2}(\mathbb{Z}_{4},U(1))=\mathbb{Z}_{1}$).
Although a priori one may not expect the topological data defining a quantum double $\mathcal{D}(G)$ to be sufficient to fully constrain the dynamics of a $G$-symmetric gapless theory, it turns out that certain boundaries of the former do encode the universal, kinematic data of the emergent anomaly discussed above. However, this requires relaxing the gapped boundary condition enforced on the dynamical boundary in Section II and considering gapless boundaries of $\mathcal{D}(G)$, which permit (partial) characterization in terms of the algebraic data governing anyon condensation. Lifting the gapped boundary condition and allowing gapless edges of the bulk TO is achieved by relaxing the constraint that the corresponding subalgebras of condensed bulk anyons $\mathcal{A}$ are Lagrangian. As illustrated in Fig. 1(a), condensing a condensable but non-Lagrangian subalgebra of anyons $\mathcal{A}$ within a subregion of the bulk TO leads to a new TO $\mathcal{D}(G)/\mathcal{A}$ within the subregion, with gapped domain walls separating the two distinct anyon theories. On the other hand, condensing a non-Lagrangian $\mathcal{A}$ at a boundary leads to a gapless boundary theory (see Fig. 1(c)). As we will review below, the reduced $\mathcal{D}(G)/\mathcal{A}$ TO encodes the effective low-energy symmetry group $G_{low}$ of the corresponding 1+11+1d gapless theory.
@@ III.1 Classification of Condensable Subalgebras and Partial Anyon Condensation
In Sec. II, we reviewed the correspondence between gapped boundaries of $\mathcal{D}(G)$ topological orders, Lagrangian subalgebras, and 1+11+1d $G$-symmetric gapped phases. Here, we review the classification of partial anyon condensation in 2+12+1d anyon theories described by the quantum double $\mathcal{D}(G)$; such partial condensations mathematically correspond to condensable (but not Lagrangian) subalgebras $\mathcal{A}$ of $\mathcal{D}(G)$, i.e., sets of bosonic anyons which all braid trivially with each other. Relaxing the Lagrangian condition amounts to lifting the condition that every anyon $b\notin\mathcal{A}$ braid nontrivially with at least one anyon $a\in\mathcal{A}$. In this scenario, anyons which lie outside $\mathcal{A}$ but braid trivially with anyons in $\mathcal{A}$ will remain deconfined anyons after $\mathcal{A}$ is condensed.
@@
The holographic dictionary described in Ref. [37] showed how this data can be understood as describing the symmetry fractionalization of $A$ and $H$ on the boundary. In the case when $H=G$, the condensable algebra $\mathcal{A}$ contains no pure charges, and the resulting boundary does not break any symmetry; since we are interested in $G$-symmetric 1+11+1d phases, we restrict to this case for the remainder of this Section. Note that the condensable algebra is Lagrangian if and only if $A=H$. In this case, $\epsilon$ is fully determined by $\eta$, and the classification reduces to the case discussed in Sec. II.
@@
where $a$ labels the condensed anyons and the integer $n_{a}\in\mathbb{Z}_{>0}$ labels the corresponding fusion multiplicity. However, this decomposition does not fully characterize the condensable algebra [88, 89]. Nevertheless, this decomposition into condensed anyons is useful in developing an intuitive picture for the condensation. Since the precise relation between the data $A\subset H$, $\epsilon$, and $\eta$ with the set of condensed anyons is somewhat involved for general $G$, we defer a detailed discussion to Appendix B and discuss the simpler case when $G$ is Abelian here.
@@
In this case, the data encoded in $\epsilon$ and $\eta$ can be repackaged into a (not necessarily Lagrangian) subgroup of condensed anyons. First, note that the condition Eq. (21b) shows that $\epsilon(a,\cdot):H\to U(1)$ is a 11d irrep of $H$. This gives the charge of the condensed dyon with flux $a$: for every flux $a\in A$, the dyon labeled by $(a,\epsilon(a,\cdot))$ is condensed. The condition Eq. (21a) ensures that the condensed anyons form a group: the charge bound to the flux $ab$ is equal to the product of the charges associated to $a$ and $b$. To understand the final condition Eq. (21c), note that since $\eta$ is a 2-cocycle on $A$, the ratio $\Omega(a,b)=\eta(b,a)/\eta(a,b)$ satisfies the same properties in Eq. (4) as the function $\Omega$ defined in Eq. (3). As before, $\Omega(a,\cdot)$ can be interpreted as the $A$-charge associated to the flux $a$. Thus, the final condition is a consistency condition that states that the restriction of the $H$-charge $\epsilon(a,\cdot)$ agrees with the $A$-charge $\Omega(a,\cdot)$ when $H$ is restricted to $A$. Finally, we note that the third condition Eq. (21c) implies that
@@ III.2 SymTFT for igSPTs
The key insight of Ref. [37] was that the data classifying condensable subalgebras of a quantum double $\mathcal{D}(G)$ is in one-to-one correspondence with the data characterizing the symmetry fractionalization patterns of $G$-symmetric igSPTs in 1+11+1d. Here, we briefly sketch out this correspondence and refer the reader to Ref. [37] for a more detailed exposition.
Consider a $G$-symmetric 1+11+1d spin system. Let us assume that, at low energies, a normal subgroup $A\lhd G$ acts trivially, such that the effective low-energy symmetry group is given by the quotient $G_{low}=G/A$. Now, on a system with periodic boundary conditions, we expect that the symmetry operator $U_{a}$, with $a\in A$, acts trivially. However, since $A$ only acts nontrivially on the gapped degrees of freedom, the symmetry may fractionalize on a system with open boundary conditions, such that $U_{a}=U_{a}^{L}\otimes U_{a}^{R}$, where $U_{a}^{L/R}$ has spatial support only on the left/right edge of the system. Since the symmetry operators $U_{a}$ form a linear representation of $A$, the restricted symmetry operators $U_{a}^{\alpha=L/R}$ need only form a projective representation:
@@
where $\eta\in\mathcal{Z}(A,U(1))$ is a 2-cocycle. Physically, the presence of a nontrivial $\eta]$ implies that the gapped degrees of freedom form an $A$-SPT.
@@
where $\epsilon(a,g)\in U(1)$ is a phase that may be interpreted as measuring the fractionalized charge under $G$ that $U^{\alpha}_{a}$ carries. As shown in Ref. [37], the 3-cocycle $\alpha]\in H^{3}(A,U(1))$ characterizing the emergent anomaly of the igSPT can be extracted from the pair of functions $(\eta,\epsilon)$.
@@ III.3 Example: $\mathbb{Z}_{4}$ igSPT
To illustrate the main idea behind the SymTFT perspective on igSPTs, let us consider in detail the example of the 1+11+1d $\mathbb{Z}_{4}$ gapless SPT first introduced in Ref. [57]. The $\mathbb{Z}_{4}$ gapless SPT is characterized by the short exact sequence
@@
where the low-energy symmetry $G_{low}=\mathbb{Z}_{2}$ exhibits an emergent anomaly characterized by the nontrivial 3-cocycle class $\alpha\in H^{3}(\mathbb{Z}_{2},U(1))\cong\mathbb{Z}_{2}$, which is trivialized by the extension to the full symmetry group $G=\mathbb{Z}_{4}$ [57]. The low energy behavior of the $\mathbb{Z}_{4}$ gapless SPT is identical to that of the symmetry-preserving boundary of the 2+12+1d $\mathbb{Z}_{2}$ Levin-Gu SPT [99], which carries the same anomaly $\alpha$. The physical consequence of this anomaly is that it prevents a 1+11+1d theory from realizing a short-range entangled state that is symmetry preserving, thereby enforcing either gaplessness or spontaneous symmetry breaking.
@@
That the data encapsulating the universal, low-energy behavior of the $\mathbb{Z}_{4}$ gapless SPT and the boundary of the Levin-Gu SPT are identical suggests that the holographic dual for the former should be the gauged version of the Levin-Gu SPT, which is given by the doubled semion model TO $\mathcal{D}(\mathbb{Z}^{\alpha}_{2})$. On the other hand, since the global symmetry for the gapless SPT is $\mathbb{Z}_{4}$, one might expect that the corresponding bulk anyon theory is given by the $\mathcal{D}(\mathbb{Z}_{4})$ TO. As it turns out, the resolution to this ostensible puzzle is that both are true since the doubled semion TO can be obtained via the $\mathcal{D}(\mathbb{Z}_{4})$ anyon theory by condensing the $e^{2}m^{2}$ anyon in the latter. The physical picture is provided by Fig. 3: the blue region denotes the $\mathcal{D}(\mathbb{Z}_{4})$ TO while the green region denotes the doubled semion TO, obtained via condensation of $e^{2}m^{2}$ in the former. Thus, the condensable algebra generated by $\langle e^{2}m^{2}\rangle$ in $\mathcal{D}(\mathbb{Z}_{4})$ precisely encodes the local low-energy properties of a 1+11+1d $\mathbb{Z}_{4}$ symmetric gapless theory (red dotted line in Fig. 3), which correspond to a $\mathbb{Z}_{2}$ symmetry along with the Levin-Gu anomaly $\alpha$. As discussed above, this is precisely the data that characterizes the $\mathbb{Z}_{4}$ igSPT.